Translate to an equation:
d dimes and q quarters total $2.00
Graph the equation and use the graph to determine three different combinations of dimes and quarters that total $2.00.

All that I have come up with is d+q=$2.00. After that I am not sure what to do.

Answer by Alan3354(69443)

Translate to an equation:
d dimes and q quarters total $2.00
Graph the equation and use the graph to determine three different combinations of dimes and quarters that total $2.00.

All that I have come up with is d+q=$2.00. After that I am not sure what to do.
--------------
d + q is the # of coins
10d + 25q = 200 Where d & q are integers
--------
2d + 5q = 40
5q = 40 - 2d
q = 8 - (2/5)d
For q to be an integer, d has t be a multiple of 5.
d....q
0....8
5....6
10...4
15...2
20...0